OpenStudy (anonymous):

I literally need your help. Find the coordinates of the stationary points on the graph of y=x^3-12x-12 and sketch the graph. Find the set of values of k for which the equation x^3-12x-12=k has more than one real solution.

OpenStudy (anonymous):

@robtobey

OpenStudy (crashonce):

stationary point means vertex

OpenStudy (crashonce):

for the second part, for there to be more than one real solution, the discriminant must be GREATER than 0

OpenStudy (crashonce):

so solve the discriminant andfind the value for where the discriminant is more than zero and it should be correct

OpenStudy (anonymous):

@CrashOnce I have a difficulty in sketching a cubic graph.?

OpenStudy (crashonce):

www.analyzemath.com/Graphing/graphing_cubic_function.html

OpenStudy (crashonce):

some help even tho im not that good lol

OpenStudy (anonymous):

lol , well I will check that out. Thanks for your help anyway :)

OpenStudy (anonymous):

same as the vertex

OpenStudy (anonymous):

ooh nvm it is not a quadratic

OpenStudy (anonymous):

\[f y=x^3-12x-12\] take the derivative \[y'=3x^2-12\] set it equal to zer0 \[3x^2-12=0\] and solve for\(x\) \[3x^2-12=0\\ x^2-4=0\\x=\pm2\]

OpenStudy (anonymous):

the first coordinate of the stationary point is \(2\) and also \(-2\) to find the second coordinate, replace \(x\) by \(2\) and \(-2\)

OpenStudy (anonymous):

you got that?

OpenStudy (anonymous):

Yes, How about skteching :) ?

OpenStudy (anonymous):

it is a cubic polynomial with leading coefficient and two stationary points gotta look something like this

OpenStudy (anonymous):

here is a better picture ' http://www.wolframalpha.com/input/?i=x^3-12x-12+

OpenStudy (jmark):

For graph http://www.wolframalpha.com/input/?i=graph++y%3Dx%5E3-12x-12 x^3-12x-12=k take the derivative k'=3x^2-12 3x^2-12=0 3x^2=12 x^2=4 x=±2 CBSE Subjects for Class 11 Arts - http://scoremore.beep.com/the-cbse-board-2014-08-13.htm